OpenStudy (anonymous):

jose sees a shirt he want to buy. His friend reminds him tha he'll have to pay the price of the shirt puls sales tax A formula describing this scenario is t=c+cb where t= total price Jose paid for the shirt c=cost of the shirt b=sales tax (percent) jose wants to find the cost of sales tax. Rewrite the formula to slove for the percent sales tax

OpenStudy (anonymous):

you need to find out b, which is the percent of sales tax. then t=c+cb or, c(1+b)=t or, (1+b)=t/c or, b=(t/c)-1

OpenStudy (anonymous):

it dosent say the percent its only said to rewrite the formula can you help me\